# Efficient Prior Selection in Gaussian Process Bandits with Thompson Sampling

### v2

@@ 4.2 Analysis of HP-GP-TS
We analyze the regret of HP-GP-TS in two ways: using the UCB-based framework of Russo & Van Roy (2014) and the information-theoretic framework of Russo & Van Roy (2016). First, note that HP-GP-TS inherits the probability matching property of GP-TS that $x_{t}|H_{t}\overset{d}{=}x^{*}|H_{t}$ where $\overset{d}{=}$ denotes equal in distribution. In addition, $p_{t}|H_{t}\overset{d}{=}p^{*}|H_{t}$ since $p_{t}$ is sampled from the posterior distribution of $p^{*}$. In the UCB-based framework, we decompose the regret into three terms:
${\mathbb{E}}[f(x^{*})-f(x_{t})]={\mathbb{E}}\Big[\underbrace{f(x^{*})-U_{t,p^{*}}(x^{*})}_{(1)}+\underbrace{U_{t,p^{*}}(x^{*})-U_{t,p^{*}}(x_{t})}_{(2)}+\underbrace{U_{t,p^{*}}(x_{t})-f(x_{t})}_{(3)}\Big]$ (8)
where $U_{t,p}(x)=\mu_{t,p}(x)+\sqrt{\beta_{t}}\sigma_{t,p}(x)$. Summing over $t\in[T]$, the first term can be bounded by a constant whilst the third term is bounded by $\sqrt{CT\beta_{T}\bar{\gamma}_{T}}$. Together, these two terms match the Bayesian regret bound for GP-TS with known prior. For the second term, one can utilize that $p^{*},x_{t}|H_{t}\overset{d}{=}p_{t},x^{*}|H_{t}$ to re-express it as ${\mathbb{E}}[U_{t,p^{*}}(x^{*})-U_{t,p_{t}}(x^{*})]$. Hence, the second term can be seen as the cost of learning the true prior. Intuitively, if the priors are similar then the upper-confidence bounds of the true and selected prior will not differ significantly. Similarly, if the priors are sufficiently distinguishable, one would expect $P_{t}(p^{*})$ to increase quickly from a few samples. In the following lemma, we demonstrate the latter for $|P|=2$ with shared kernel functions. {lemma}[] If $|P|=2$ and the two priors share the same kernel function ($k_{p}=k$ $\forall p\in P$), then for any fixed sequence of arms $x_{1:t}=\{x_{i}\}_{i=1}^{t}$ the posterior probability of the true prior $p^{*}$ satisfies
${\mathbb{E}}_{{\mathbf{y}}}\left[P_{t+1}(p)\big|p^{*}=p,x_{1:t}\right]\geq 1+P_{0}(p)e^{\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert^{2}}\Phi\Big(-\frac{3}{2}\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert\Big)-\frac{1}{P_{0}(p)}\Phi\Big(-\frac{\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert}{2}\Big)$ (9)
where ${\bm{\mu}}\in{\mathbb{R}}^{t}$ such that $({\bm{\mu}})_{i}=\mu_{p^{*}}(x_{i})-\mu_{\tilde{p}}(x_{i})$ for $\tilde{p}\neq p^{*}$ and $({\bm{\Sigma}})_{i,j}=k(x_{i},x_{j})$. From Eq.˜8, we note that $P_{t}(p^{*})$ increases the most when $\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert$ is maximized, which corresponds to selecting arms with low correlation and large difference in prior mean (relative to the variance). Hence, $P_{t}(p^{*})$ increases quickly at first when HP-GP-TS explores but levels out when the algorithm starts exploiting. Next, we state the Bayesian regret bound for HP-GP-TS.
@@
${\text{BR}}(T)\leq\pi^{2}/6+{\sum_{t\in[T]}}{\mathbb{E}}\left[U_{t,p^{*}}(x^{*})-U_{t,p_{t}}(x^{*})\right]+\sqrt{CT\beta_{T}\bar{\gamma}_{T}}.$ (10)
@@
Unlike PE-GP-TS and PE-GP-UCB, the regret bound of the third term for HP-GP-TS depends on the average MIG $\sqrt{\bar{\gamma}_{T}}$ rather than the worst case $\sqrt{|P|\hat{\gamma}_{T}}$ which can impact the theoretical regret significantly if the complexity of the priors differ and the prior is weighted towards simple priors. This is reasonable since the elimination methods assume arbitrary selection of $p^{*}$ as opposed to sampling from a hyperprior. If the hyperprior is deterministic then the regret bound for HP-GP-TS matches that of GP-TS up to a factor ${\mathcal{O}}(\sqrt{\log T})$ (Takeno et al., 2024) and $\bar{\gamma}_{T}$ would be equal to the worst case $\hat{\gamma}_{T}$. Again, using the lower bound of Scarlett (2018), our upper bound would be tight up to a factor of ${\mathcal{O}}(\sqrt{\beta_{T}\bar{\gamma}_{T}})$.
The information-theoretic framework of Russo & Van Roy (2016) can be applied generally if the probability matching property is satisfied and the rewards are subgaussian (Vashishtha & Maillard, 2025).
Theorem 4.3.
If $p^{*}\sim P_{1}$, $f\sim{\mathcal{GP}}(\mu_{1,p^{*}},k_{1,p^{*}})$, the Bayesian regret of HP-GP-TS is bounded by
${\text{BR}}(T)\leq\sqrt{2|{\mathcal{X}}|\log(|{\mathcal{X}}|)(\sigma^{2}_{0}+\sigma^{2})T}.$ (11)
The proof of Theorem˜4.3 follows the proof in section D.2 of Russo & Van Roy (2016) with subgaussian noise. For completeness, we provide a proof using the GP-bandit notation in Algorithm˜2. The information-theoretic regret bound is ${\mathcal{O}}(\sqrt{T})$ which improves upon the ${\mathcal{O}}(\sqrt{T\beta_{T}\bar{\gamma}_{T}})$ obtained previously and would match the lower bound of Scarlett (2018) in terms of $T$. However, we also obtain a ${\mathcal{O}}(\sqrt{|{\mathcal{X}}|\log|{\mathcal{X}}|})$ dependency.
@@
Figure 3: Cumulative regret for synthetic experiments with varying kernel (left), lengthscale (center) and mean function (right). The average final regret for PE-GP-UCB is 116.5 and 389.0 in the lengthscale and subspace experiments. Errorbars correspond to $\pm 1$ standard error.
@@
5 Experiments
Synthetic experiments
We consider three synthetic setups with different choices of priors in $P$. For the first setup, the priors have one of the following kernels: i) RBF kernel, ii) the rational quadratic kernel with $\alpha=0.5$, iii) Matérn kernel with $\nu=5/2$, iv) Matérn kernel with $\nu=3/2$, v) periodic kernel with period $\rho=5$, vi) linear kernel with $v=0.05^{2}$. All kernels use a lengthscale of $1.0$ and are scaled s.t. $k(x,\tilde{x})\leq 1$. In addition, the mean function for all priors is zero everywhere. For the second setup, the priors use the RBF kernel with lengthscales $4$, 22, 11 or $1/2$. For the third setup, the total dimensions $d=16$ but each prior $p_{i}$ assumes $f(x)$ depends on $d_{s}=4$ subdimensions: $[i,i+1,i+2,i+3]$ for $i\in[5]$. Dimensions larger than $5$ are wrapped around 1, i.e. $((j-1)\mod 5)+1$, such that the priors are equally difficult to distinguish and optimize. All priors use the RBF kernel with lengthscale $\ell=8$. For all three setups, the true prior $p^{*}$ is sampled uniformly from $P$, the noise variance $\sigma^{2}=0.25^{2}$, and the horizon $T=500$. For the first two setups, 500 arms are equidistantly spaced in $[0,20]$ and for the third 500 arms are sampled uniformly on $0,20]^{16}$. The prior elimination methods use $\delta=0.05$. All models are evaluated on 500 seeds on each setup. As baselines, we use PE-GP-UCB and Maximum A Posteriori (MAP) GP-TS where MAP GP-TS is identical to HP-GP-TS except for greedily selecting $p_{t}$ from the posterior: $p_{t}=\operatorname*{arg\,max}_{p}P_{t-1}(p)$ [^1] Note that since the hyperprior is uniform, MAP is equivalent to discrete maximum likelihood estimation.. Regardless of the selected prior, the weighting of the hyperposterior $P_{t}$ is updated for all priors. Hence, greedily selecting the prior could reduce unnecessary exploration. In addition, we investigate the oracle variants of PE-GP-TS and PE-GP-UCB with $\delta=0.05$ that are only given the true prior: $P_{1}=\{p^{*}\}$.
The cumulative regret for the three synthetic experiments is shown in Fig.˜3. Across all three experiments, we observe that HP-GP-TS has lower regret than the other methods and performs close to the oracle GP-TS. For the kernel and subspace experiments, PE-GP-TS has lower regret than the oracle GP-UCB. Hence, even if PE-GP-UCB was optimized to perform as well as the oracle, it would still not achieve the regret of our proposed methods. MAP GP-TS has slightly higher regret than HP-GP-TS for the lengthscale and subspace experiments but has significantly higher regret and variance for the kernel experiment. The greedy selection of MAP (MLE) leads to under-exploration for MAP GP-TS in certain instances.
The number of priors remaining $|P_{t}|$ and the hyperposterior entropy for the kernel experiment is shown in Fig.˜4. The PE-methods eliminate at most one prior on average. In contrast, the hyperposterior entropy of HP-GP-TS is equivalent to 80-90% of the probability mass being assigned to one prior. HP- and MAP-GP-TS thus effectively discards priors at a much faster rate. The same pattern holds for the lengthscale and subspace experiments, see Figs.˜9 and 10 in Appendix˜D.
@@ 5 Experiments
In Fig.˜5, we visualize how often the methods select the true prior $p^{*}$ (or kernel) in the kernel experiment as confusion matrices. PE-GP-UCB selects the Matérn-3/2 kernel more than 96% of the rounds. The Matérn-3/2 kernel induces a distribution over functions that are less smooth compared to the other kernels and produces much higher confidence intervals outside the observed data leading to excessive optimistic exploration. PE-GP-TS also shows a bias towards the Matérn-3/2 kernel but does not select it as frequently as PE-GP-UCB - demonstrating that one layer of optimism has been removed. The overall “accuracy" of the selected priors, i.e. ${\sum_{t\in[T]}}\mathds{1}\{p_{t}=p^{*}\}/T$, for the elimination-based methods is around $17\%$ in the kernel experiment compared to $62.5\%$ and $63.2\%$ for MAP and HP-GP-TS respectively. For HP-GP-TS, we observe that it can easily identify the periodic and linear kernels. However, the RBF, Matérn and RQ kernels are often confused with each other. These kernels do not have as easily distinguishable characteristics and are likely to produce similar posteriors even with a small amount of data. See Fig.˜11 in Appendix˜D for confusion matrices in the lengthscale and subspace experiments.
Figure 6: Total regret for the lengthscale and subspace experiments as the number of priors $|P|$ increases.
@@
We perform two experiments to understand how the regret of our algorithms scale with the number of priors. In both experiments, the average difficulty of the problem is kept constant such that the regret of the oracle models is constant. In the first experiment, we increase the discretization of the lengthscale values. The lengthscales are equidistantly spaced in $[0.5,4]$ with $|P|\in\{8,16,32,64,128\}$. As $|P|$ increases, the difference between similar priors is reduced. In the second experiment, we increase the number of priors in the subspace experiment from 5 up to 16. Each prior can share at most 3 out of 4 dimensions with other priors which ensures the priors remain meaningfully different. The total regret as the number of priors increases is shown in Fig.˜6. For the lengthscale experiment, increasing the number of priors above 8 does not affect the regret for any algorithm. This is likely due to the redundancy in the priors. However in the subspace experiment, the regret of the prior elimination algorithms scales approximately as $\sqrt{|P|}$ whilst MAP- and HP-GP-TS are consistently close to the constant regret of the oracle.
Figure 7: Cumulative regret on Intel temperature data (top) and PeMS data (bottom). Errorbars correspond to $\pm 1$ standard error. The average final regret for PE-GP-UCB is 511.9 for PNW.
@@
We perform three experiments with real-world data from the Intel Berkeley dataset (Madden et al., 2004), California Performance Measurement System (PeMS) (Chen et al., 2001; California Department of Transportation, 2024) and Pacific Northwest (PNW) daily precipitation dataset (Widmann & Bretherton, 1999, 2000). Each dataset contains measurements from a set of sensors over time. We split each dataset into a training and test set where the test set contains the last third of the data. Hence, the distribution of the test data may have shifted slightly from the training data. The training sets are split further into separate buckets to define our priors. For each bucket $p$, we compute the empirical mean $\hat{\mu}_{p}$ and covariance $\hat{\Sigma}_{p}$ which defines the prior ${\mathcal{GP}}(\hat{\mu}_{p},\hat{\Sigma}_{p})$. The buckets in the Intel data corresponds to the 12 days in the training dataset. For the PeMS data, each hour between 06:00 and 13:00 defines one prior, giving 7 priors. For the daily precipication data, each month in the year constitutes a prior, yielding 12 priors. When running the experiments, we select a measurement of all sensors from the test data uniformly at random. The selected measurements correspond to the unknown function $f(x)$ where $x$ is the sensor index and the goal is then to identify sensors measuring large temperatures, small speeds or high precipication respectively for the three datasets. When the algorithms select an arm to evaluate, we add Gaussian noise to $y_{t}$ with variance $\sigma^{2}$ around 5% of the signal variance, similar to Srinivas et al. (2012); Bogunovic et al. (2016). See Appendix˜C for more details about the experimental setup.
The cumulative regret for the experiments with real-world data is presented in Fig.˜7. For the Intel and PNW data, HP- and MAP GP-TS obtain the lowest and second lowest cumulative regret respectively. HP- and MAP GP-TS have lower regret than PE-GP-TS initially but PE-GP-TS catches up and has the lowest total regret. To understand this better, we visualize quantiles of the total regret in Fig.˜13. MAP- and HP-GP-TS have the lowest median regret for all three experiments and hence perform best in a majority of instances. However, the 90th and 95th quantiles are considerably larger for the PeMS data which impacts the average regret significantly. Hence, for the PeMS data, the prior elimination methods seem to yield more stable results.
In Fig.˜4, the number of priors remaining in $|P_{t}|$ and the hyperposterior entropy is shown for the Intel experiment. Similar to the synthetic experiment, on average, the prior elimination methods eliminate less than 1 prior whereas the hyperposteriors of HP- and MAP GP-TS concentrate the equivalent of 80-90% of the probability mass to one prior. The results for PeMS and PNW experiment are shown in Figs.˜9 and 10 in Appendix˜D. Here, effectively no priors are eliminated. For the PNW experiment, the hyperposteriors do not concentrate as much compared to the other experiments. This could indicate that knowing the exact prior is not as important for the PNW data.

### v3

@@ 4.2 Analysis of HP-GP-TS
We analyze the regret of HP-GP-TS using the UCB-based framework of Russo and Van Roy (2014). First, note that HP-GP-TS inherits the probability matching property of GP-TS that $x_{t}|H_{t}\overset{d}{=}x^{*}|H_{t}$ where $\overset{d}{=}$ denotes equal in distribution. In addition, $p_{t}|H_{t}\overset{d}{=}p^{*}|H_{t}$ since $p_{t}$ is sampled from the posterior distribution of $p^{*}$. We decompose the regret into three terms:
${\mathbb{E}}[f(x^{*})-f(x_{t})]={\mathbb{E}}\Big[\underbrace{f(x^{*})-U_{t,p^{*}}(x^{*})}_{(1)}\\ +\underbrace{U_{t,p^{*}}(x^{*})-U_{t,p^{*}}(x_{t})}_{(2)}+\underbrace{U_{t,p^{*}}(x_{t})-f(x_{t})}_{(3)}\Big]$ (8)
where $U_{t,p}(x)=\mu_{t,p}(x)+\sqrt{\beta_{t}}\sigma_{t,p}(x)$. Summing over $t\in[T]$, the first term can be bounded by a constant whilst the third term is bounded by $\sqrt{CT\beta_{T}\bar{\gamma}_{T}}$. Together, these two terms match the Bayesian regret bound for GP-TS with a known prior. For standard GP-TS when the true prior $p^{*}$ is known, the second term has expectation zero. However for HP-GP-TS, one can utilize that $p^{*},x_{t}|H_{t}\overset{d}{=}p_{t},x^{*}|H_{t}$ to re-express the second term as ${\mathbb{E}}[U_{t,p^{*}}(x^{*})-U_{t,p_{t}}(x^{*})]$. Hence, the second term can be seen as the cost of learning the true prior. Intuitively, if the priors are similar then the upper-confidence bounds of the true and selected prior will not differ significantly. Similarly, if the priors are sufficiently distinguishable, one would expect $P_{t}(p^{*})$ to increase quickly from a few samples. In the following lemma, we demonstrate the latter for $|P|=2$ with shared kernel functions. {lemma}[] If $|P|=2$ and the two priors share the same kernel function ($k_{p}=k$ $\forall p\in P$), then for any fixed sequence of arms $x_{1:t}=\{x_{i}\}_{i=1}^{t}$ the posterior probability of the true prior $p^{*}$ satisfies
${\mathbb{E}}_{{\mathbf{y}}}\left[P_{t+1}(p)\big|p^{*}=p,x_{1:t}\right]\geq 1+P_{0}(p)e^{\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert^{2}}\\ \times\Phi\Big(-\frac{3}{2}\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert\Big)-\frac{1}{P_{0}(p)}\Phi\Big(-\frac{\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert}{2}\Big)$ (9)
where ${\bm{\mu}}\in{\mathbb{R}}^{t}$ such that $({\bm{\mu}})_{i}=\mu_{p^{*}}(x_{i})-\mu_{\tilde{p}}(x_{i})$ for $\tilde{p}\neq p^{*}$ and $({\bm{\Sigma}})_{i,j}=k(x_{i},x_{j})$. In Equation˜8, $\bm{\mu}$} is the difference in mean between the two priors for the given sequence of arms and ${\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}$ is the whitened difference vector. As the norm of the whitened difference vector increases, $\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert$, the sum of the two rightmost terms in Equation˜9 goes to zero from below. The norm $\lVert{\bm{\Sigma}^{-\frac{1}{2}}}{\bm{\mu}}\rVert$, and consequently the posterior probability of the true prior $P_{t+1}(p^{*})$, is maximized when arms with large difference in prior mean (relative to their variance) and with low correlation are selected. Hence, $P_{t}(p^{*})$ increases quickly at first when HP-GP-TS explores but levels out when the algorithm starts exploiting.
Next, we state the Bayesian regret bound for HP-GP-TS.
@@
${\text{BR}}(T)\leq\pi^{2}/6+{\sum_{t\in[T]}}{\mathbb{E}}\left[U_{t,p^{*}}(x^{*})-U_{t,p_{t}}(x^{*})\right]\\ +\sqrt{CT\beta_{T}\bar{\gamma}_{T}}.$ (10)
@@
Unlike PE-GP-TS and PE-GP-UCB, the regret bound of the third term for HP-GP-TS depends on the average MIG $\sqrt{\bar{\gamma}_{T}}$ rather than the worst case $\sqrt{|P|\hat{\gamma}_{T}}$ which can impact the theoretical regret significantly if the complexity of the priors differ and the prior is weighted towards simple priors. This is reasonable since the elimination methods assume arbitrary selection of $p^{*}$ as opposed to sampling from a hyperprior. If the hyperprior is deterministic then the regret bound for HP-GP-TS matches that of GP-TS up to a factor ${\mathcal{O}}(\sqrt{\log T})$ (Takeno et al., 2024) and $\bar{\gamma}_{T}$ would be equal to the worst case $\hat{\gamma}_{T}$. Again, using the lower bound of Scarlett (2018), our upper bound would be tight up to a factor of ${\mathcal{O}}(\sqrt{\beta_{T}\bar{\gamma}_{T}})$ when considering only the first and third terms. We have not shown that ${\sum_{t\in[T]}}{\mathbb{E}}[U_{t,p^{*}}(x^{*})-U_{t,p_{t}}(x^{*})]$ is sublinear. As noted before, this sum corresponds to the cost of learning the true prior. Conceptually, we expect the terms in the sum to decrease over time as either $p_{t}$ or $x_{t}$ converge to $p^{*}$ or $x^{*}$ respectively. Empirically, we observe this in our synthetic experiments, see Figure˜13 in Appendix˜F.
4.3 Comparison to MixTS
Hong et al. (2022a) study MixTS, a Thompson sampling algorithm that assumes the prior is a mixture distribution, for standard and linear bandits. Using our GP-notation, their analysis starts from the decomposition
${\mathbb{E}}[f(x^{*})-f(x_{t})]={\mathbb{E}}[\underbrace{f(x^{*})-\mu_{t,p^{*}}(x^{*})}_{(1)}\\ +\underbrace{\mu_{t,p_{t}}(x_{t})-f(x_{t})}_{(2)}].$ (11)
The first term is bounded using standard UCB techniques. For the second term, Hong et al. (2022a) introduce confidence sets for the priors $C_{t}$ based on the excess reward each prior observes and prove that $p^{*}\in C_{t}$ with high probability. Their key idea is to introduce confidence sets for the priors $C_{t}$ based on the excess reward each prior observes and prove that $p^{*}\in C_{t}$ holds with high probability, similar to Ziomek et al. (2025). In our GP setting, their excess reward can be defined as
$G_{t}(p)=\sum_{s=1}^{t-1}\mathds{1}\{p_{s}=p\}(\mu_{s,p}(a_{s})-\sqrt{\beta_{t}}\sigma_{s,p}(a_{s})-y_{s})$ (12)
and the prior confidence set $C_{t}=\{p\in P:G_{t}(p)\leq 2\sigma\sqrt{N_{t}(p)\log T}\}$ where $N_{t}(p)$ is the number of times prior $p$ has been selected before time $t$. Unlike the definition of Ziomek et al. (2025), priors can return to the confidence set after previously being removed from it. When bounding the second term in their decomposition for a standard bandit with bounded rewards, Hong et al. (2022a) condition on $p_{t}\in C_{t}$. If $p_{t}\in C_{t}$, the second term can be bounded by all the excess $\sum_{p\in P}G_{T+1}(p)$. If $p_{t}\notin C_{t}$ and the rewards are in $[0,1]$, then the second term is bounded by the probability of $p^{*}\in C_{t}$. For the linear setting with unbounded rewards, the proof requires conditioning on the linear parameter vector $\theta^{*}$ to lie close to its prior mean. Under this additional event $E_{0}$, the distribution of the true parameter vector $\theta^{*}$, the true prior and the optimal arm can shift. But, conditioned on the history $H_{t}$, MixTS is unaffected by $E_{0}$ at time step $t$. However, Hong et al. (2022a) use that $x^{*},p^{*}|H_{t},E_{0}\overset{d}{=}x_{t},p_{t}|H_{t},E_{0}$ without proof, potentially invalidating Theorem 1 of Hong et al. (2022a). We describe this and other technical issues in detail in Appendix˜B.
@@ 4.3 Comparison to MixTS
Figure 3: Cumulative regret for synthetic experiments with varying kernel (left), lengthscale (center) and mean function (right). The final regret for PE-GP-UCB is 114 and 389 in the lengthscale and subspace experiments, and 181 for SCoreBO in the lengthscale experiment. Errorbars correspond to $\pm 1$ standard error.
5 Experiments
In this section, we describe our experiments based on synthetic and real-world data.
Synthetic experiments
We consider three synthetic setups with different choices of priors in $P$. For the first setup, the priors have one of the following kernels: i) RBF kernel, ii) the rational quadratic kernel with $\alpha=0.5$, iii) Matérn kernel with $\nu=5/2$, iv) Matérn kernel with $\nu=3/2$, v) periodic kernel with period $\rho=5$, vi) linear kernel with $v=0.05^{2}$. For the second setup, 8 priors use the RBF kernel with different lengthscales equidistantly spaced between $1/2$ and $4$. For the third setup, the total dimensions $d=16$ but each of the 5 priors $p_{i}$ assumes $f(x)$ depends on $d_{s}=4$ subdimensions. The 4 subdimensions are designed such that the priors are equally difficult to distinguish. All priors use the RBF kernel with lengthscale $\ell=8$. For all three setups, the true prior $p^{*}$ is sampled uniformly from $P$, the noise variance $\sigma^{2}=0.25^{2}$, and the horizon $T=500$. For the first two setups, 500 arms are equidistantly spaced in $[0,20]$ and for the third 500 arms are sampled uniformly on $0,20]^{16}$. All models are evaluated on 500 seeds on each setup. As baselines, we use PE-GP-UCB, SCoreBO, fully Bayesian Expected Improvement (EEI) and Maximum A Posteriori (MAP) GP-TS. MAP GP-TS is identical to HP-GP-TS except for greedily selecting $p_{t}$ from the posterior: $p_{t}=\operatorname*{arg\,max}_{p}P_{t-1}(p)$ [^1] Note that since the hyperprior is uniform, MAP is equivalent to discrete maximum likelihood estimation.. In addition, we compare against the oracle variants of PE-GP-TS and PE-GP-UCB that are only given the true prior: $P_{1}=\{p^{*}\}$.
The cumulative regret for the three synthetic experiments is shown in Figure˜3 and the final regret is shown in Table˜1 in Appendix˜F. Across all three experiments, we observe that HP-GP-TS and EEI has lower regret than the other methods and performs close to the oracle GP-TS. For the kernel and subspace experiments, PE-GP-TS has lower regret than the oracle GP-UCB. Hence, even if PE-GP-UCB was optimized to perform as well as the oracle, it would still not achieve the regret of our proposed methods. MAP GP-TS has slightly higher regret than HP-GP-TS for the lengthscale and subspace experiments but has significantly higher regret and variance for the kernel experiment. The greedy selection of MAP (MLE) leads to under-exploration for MAP GP-TS in certain instances. SCoreBO has the highest regret in the kernel and lengthscale experiment but has more comparable performance in the subspace experiment. In Figure˜10 in Appendix˜F, we report the regret for the two most competitive methods, HP-GP-TS and EEI, with an extended horizon $T=1500$ where we observe that HP-GP-TS yields noticeably lower regret.
The number of priors remaining $|P_{t}|$ and the hyperposterior entropy for the kernel experiment is shown in Figure˜4. The PE-methods eliminate at most one prior on average. In contrast, the final hyperposterior entropy across all algorithms is equivalent to 70-99% of the probability mass being assigned to one prior showing that the hyperposterior adapts more effectively. Across the experiments, SCoreBO has the lowest hyperposterior entropy followed by HP-GP-TS and EEI has the highest (except for the kernel experiment), see Figures˜11 and 12 in Appendix˜F. Thus, HP-GP-TS has similar regret to EEI but lower hyperposterior entropy.
In Figure˜5, we visualize how often the methods select the true prior $p^{*}$ (or kernel) in the kernel experiment as confusion matrices. PE-GP-UCB selects the Matérn-3/2 kernel more than 96% of the rounds. The Matérn-3/2 kernel induces a distribution over functions that are less smooth compared to the other kernels and produces much higher confidence intervals outside the observed data leading to excessive optimistic exploration. PE-GP-TS also shows a bias towards the Matérn-3/2 kernel but does not select it as frequently as PE-GP-UCB - demonstrating that one layer of optimism has been removed. The overall “accuracy" of the selected priors, i.e. ${\sum_{t\in[T]}}\mathds{1}\{p_{t}=p^{*}\}/T$, for the elimination-based methods is around $17\%$ in the kernel experiment compared to $62.9\%$ and $63.2\%$ for MAP and HP-GP-TS respectively. For HP-GP-TS, we observe that it can easily identify the periodic and linear kernels. However, the RBF, Matérn and RQ kernels are often confused with each other. These kernels do not have as easily distinguishable characteristics and are likely to produce similar posteriors even with a small amount of data. See Figure˜14 in Appendix˜F for confusion matrices in the lengthscale and subspace experiments.
@@ 5 Experiments
Figure 6: Total regret for the lengthscale and subspace experiments as $|P|$ increases.
Figure 7: Cumulative regret on the real-world data experiments. Errorbars correspond to $\pm 1$ standard error. The average final regret for SCoreBO and PE-GP-UCB is 861 and 506 on PNW.
@@
We perform two experiments to understand how the regret of our algorithms scale with the number of priors. In both experiments, the average difficulty of the problem is kept constant such that the regret of the oracle models is constant. In the first experiment, we increase the discretization of the lengthscale values. The lengthscales are equidistantly spaced in $[0.5,4]$ with $|P|\in\{8,16,32,64,128\}$. As $|P|$ increases, the difference between similar priors is reduced. In the second experiment, we increase the number of priors in the subspace experiment from 5 up to 16. Each prior can share at most 3 out of 4 dimensions with other priors which ensures the priors remain meaningfully different. The total regret as the number of priors increases is shown in Figure˜6. For the lengthscale experiment, increasing the number of priors above 8 does not affect the regret for any algorithm, likely due to the increased redundancy in the priors. The one exception is SCoreBO, whose regret increases as $|P|$ is increased to 32 but levels off beyond that. In the subspace experiment, the regret of the prior elimination algorithms scales approximately as $\sqrt{|P|}$ whilst MAP- and HP-GP-TS are consistently close to the constant regret of the oracle. The regret of EEI drops initially but is otherwise constant and SCoreBO also has constant regret.
@@
We perform three experiments with real-world data from the Intel Berkeley dataset (Madden and others, 2004), California Performance Measurement System (PeMS) (Chen et al., 2001; California Department of Transportation, 2024) and Pacific Northwest (PNW) daily precipitation dataset (Widmann and Bretherton, 1999, 2000). Each dataset contains measurements from a set of sensors over time. We split each dataset into a training and test set where the test set contains the last third of the data. Hence, the distribution of the test data may have shifted slightly from the training data. The training sets are split further into separate buckets to define our priors. For each bucket $p$, we compute the empirical mean $\hat{\mu}_{p}$ and covariance $\hat{\Sigma}_{p}$ which defines the prior ${\mathcal{GP}}(\hat{\mu}_{p},\hat{\Sigma}_{p})$. The buckets in the Intel data corresponds to the 12 days in the training dataset. For the PeMS data, each hour between 06:00 and 13:00 defines one bucket, giving 7 priors. For the daily precipitation data, each month in the year constitutes a bucket, yielding 12 priors. When running the experiments, we select a measurement of all sensors from the test data uniformly at random. The selected measurements correspond to the unknown function $f(x)$ where $x$ is the sensor index and the goal is then to identify sensors measuring large temperatures, small speeds or high precipitation respectively for the three datasets. When the algorithms select an arm to evaluate, we add Gaussian noise with variance $\sigma^{2}$ around 5% of the signal variance, similar to Srinivas et al. (2012); Bogunovic et al. (2016). See Appendix˜E for details about the experimental setup.
The cumulative regret for the experiments with real-world data is presented in Figure˜7. Across these experiments, HP-GP-TS has either the lowest regret or is within 1 standard error of the algorithm with the lowest regret. SCoreBO has significantly higher regret than all other methods. Notably for the PeMS data, PE-GP-TS has the lowest average regret whereas MAP GP-TS and EEI perform worse compared to the other experiments.
The number of priors remaining in $|P_{t}|$ and the hyperposterior entropy for the real-world data experiments is shown in Figures˜11 and 12 in Appendix˜F. Similar to the synthetic experiments, on average, the prior elimination methods eliminate less than 1 prior at best and no priors (across all 500 seeds) at worst. In contrast, the hyperposterior of HP-GP-TS concentrates to the equivalent of 60-80% of the probability mass to one prior. The relative standing in terms of reduced hyperposterior uncertainty between SCoreBO, HP-GP-TS and EEI remains consistent across all the experiments. SCoreBO reduces the hyperposterior uncertainty the most at the cost of significantly higher regret whereas HP-GP-TS provides a better balance between low regret and low hyperposterior uncertainty.